# ASTRODEEP-JWST: NIRCam-HST multiband photometry and redshifts for half a million sources in six extragalactic deep fields

### v1

@@ 6 Summary and conclusions
We presented and discussed a major release of photometric catalogs, collecting data from eight JWST observational programs on six deep extra-galactic fields. We release a new reduction of the Abell 2744 composite mosaics, gathering the GLASS-JWST, UNCOVER, DDT2756 and GO3990 programs, and use the reductions provided by the teams of the other programs. NIRCam data are complemented with archival HST images on which we perform new measurements.
The catalogs, mainly conceived for high redshift science, include a grand-total of 531,173 objects, of which 17,359 are tagged as spurious, and 2,268 more have bad photometric measurements ($\geq$400, see Section 3.5). Sources were detected with SExtractor on stacks of the F356W and F [^4] W mosaics of each field. For all the detections we provide positions in equatorial and pixel coordinates, basic morphological parameters, total fluxes and corresponding uncertainties in 16 photometric bands computed by means of the software a-phot, a diagnostic flag, and four photometric redshift estimates obtained with zphot and EAzY.
We performed validation tests on the astrometry, the photometry and the photo-$z$ accuracy, comparing our catalogs against other releases and finding a general good agreement.
The catalogs are available for download from the Astrodeep website, https://astrodeep.eu. Updates and/or additional releases will be documented on the website and uploaded to the same repository. We encourage the community to exploit these catalogs for any suitable scientific purpose.
@@ Acknowledgements.
We also acknowledge support from the INAF Large Grant 2022 “Extragalactic Surveys with JWST” (PI Pentericci).
@@
Appendix A Catalog format
We release several catalogs for each field, among which one (named “optap”-catalog) is obtained assigning to each source the colors computed in its “optimal” aperture (see Sect. 3.4). The catalogs are published in fits format, and come with a README file that explains the meaning and format of the columns. All catalogs first list a set of columns indicating the unique identifiers of the detected objects, their position in equatorial coordinates and in pixels, and some values from the SExtractor detection run on the F356W+F [^4] W stack mosaics: ISOAREA_IMAGE, CLASS_STAR, FLAGS, FLUX_RADIUS, FLUX_AUTO, FLUXERR_AUTO. Then, some measures obtained with a-phot again on the detection stack: the major semi-axis of the elliptical isophote, the ellipticity and the position angle, and the Kron radius (in pixels). The total fluxes in the 16 photometric bands, computed as described in Sect. 3.4, are then given, followed by the corresponding 16 uncertainties (all in $\mu$Jy units). Finally, the last column lists the flags discussed in Sect. 3.5. In the “optap”-catalog the value of the “optimal” aperture is also given (in arcseconds).
We separately release a set of photometric redshifts catalogs estimated on the “optap”-catalog with local background subtraction. For each source they list the ID, RA and Dec, the spectrscopic redshift when available, the four estimates of the photometric redshift (see Sect. 5), and the flag assigned to the source as reported also in the “optap”-catalog, with the additional indication of a negative sign for the sources lacking information in the bands necessary to identify the Lyman break.
@@ Appendix B Astrometry
As explained in Sect. 3.1, we re-aligned HST mosaics to NIRCam grid by means of an automatic procedure. Table 6 lists the offsets in RA and Dec after the procedure, for all the bands and fields. We point out that the CANDELS EGS and CEERS HDR1 HST data are obtained from the same observations, but the latter has been re-aligned to Gaia-DR3; since we took care to correct all of the images for astrometric accuracy, the two datasets should be perfectly consistent.
Table 6: The median and MAD of the astrometric offsets of the HST bands, computed cross-matching the sources and corrected before re-projecting the HST images to the JWST grid. For each field, the F160W values are the coordinate offsets between the original HST F160W and the NIRCam F150W band, while the others refer to the offsets between the given band and F160W band after it was re-aligned to the F150W.
Target Band $N_{obj.}$ $\Delta{RA}$ (”) $\Delta{DEC}$ (”)
ABELL2744 F435W 1307 0.009 $\pm$ 0.02 $0$$\pm$ 0.01
@@
Appendix C Selection of stars and potentially spurious sources
@@
We point out that this approach joins two similar and complementary techniques using diagnostic diagrams to identify the loci occupied by point-like and spurious sources, i.e. the SNR vs radius plane and the $\mu$-mag vs mag plane. Fig. 9 shows the PCA plane for all fields in the left column, and the two diagrams corresponding to the other techniques in the central and right columns.
@@
Figure 9: Star/galaxy separation and flagging of spurious sources. From left to right: as obtained in the PCA plane described in Sect. 3.5, projected on the hlr vs. SNR plane, projected on the $\mu$MAX-MAG vs. MAG plane. The last column shows the values of the FWHM in the detection F356W+F [^4] W stack, for the sources identified as point-source.
Appendix D Depth of mosaics
@@
[…]
Appendix E Comparisons with other photometric catalogs
Fig. 11 to 16 show how the photometry obtained in this work compares to that from other available catalogs. For each field we consider recent NIRCam-based catalogs and archival HST based data, from CANDELS or Astrodeep. For the latter, we used the $Ks$ and IRAC CH1-2 bands as proxys for the F200W, F356W and F [^4] W JWST bands, applying the following color corrections, based on Bruzual & Charlot (2003) theoretical templates: $Ks$=F200W-0.06, IRAC1=F356W+0.02, IRAC2=F [^4] W-0.01.
Each panel of the plots shows the comparison for one color, with the relative difference $\Delta c/c$ (i.e. color measured in this work minus color in the reference work, divided by color in the archival work) as a function the magnitude of the second band in this work (e.g., F [^4] W magnitude if the color is F356W-F [^4] W). The blue line is the median of the distribution. We consider sources with SNR$>$5 and flag$<$200. See Sect.4 for more details.
Figure 11: Comparison of colors measured in this work and archival catalogs. Shown are relative errors $\Delta c$ i.e. the colors measured in this work minus those in the reference catalogs), vs. the second band magnitude in this work catalog (e.g. F [^4] W for the F356-F [^4] W color). Top to bottom: ABELL2744 vs. P23 and NGDEEP vs. Finkelstein’s catalog (priv. comm.). The number of cross-matched sources after excluding those with SNR $<5$ in any of the two catalogs or flag $\geq 200$ in this work is also given in each panel; the blue line is the median of the distribution.
Figure 12: Same as Fig. 11, for (top to bottom) ABELL2744 vs. Weaver et al. (2024) and CEERS vs. Finkelstein et al. (2023a).
Figure 13: Same as Fig. 11, for JADES-GN and JADES-GS vs. the JADES team catalogs (Rieke et al. 2023).
Figure 14: Same as Fig. 11, for ABELL2744 vs. Merlin et al. (2016b) showing individual points rather than a density hexbin plot because of the small number of cross-matched sources, and PRIMER-COSMOS vs. Nayyeri et al. (2017). The HST colors are transformed into JWST colors by means of the corrections described in Appendix E.
Figure 16: Same as Fig. 11, JADES-GS vs. Merlin et al. (2021) and NGDEEP vs. Merlin et al. (2021), the latter showing individual points rather than a density hexbin plot because of the small number of cross-matched sources. The HST colors are transformed into JWST colors by means of the corrections described in Appendix E.
Appendix F Photometric redshifts validation
Table 8: Compared statistics for photometric vs. spectroscopic redshifts. The listed values are: number N of matched sources (the number can vary due to failings in the fitting procedures of the different codes); mean, standard deviation, median and NMAD of the quantity $|z_{phot}-z_{spec}|/(1+z_{spec})$, all multiplied by 100 to make the differences clearer; and percentage of outliers $f_{outliers}$ as described in Sect. 5, for each run using a different software and/or templates, and for each field.
@@ Appendix F Photometric redshifts validation
zphot 1209 0.72 4.18 0.57 3.28 9.35
EAzY v1.3 1208 -0.77 4.24 -0.37 3.25 12.83
EAzY Larson Lya 1192 0.04 4.53 -0.14 3.21 14.18
EAzY Larson LyaRed 1192 -0.12 4.58 -0.19 3.26 14.77
@@
zphot 3081 1.19 4.63 0.92 3.69 7.95
EAzY v1.3 3077 0.37 4.51 0.44 3.59 7.47
EAzY Larson Lya 3076 -0.41 4.96 -0.29 4.05 8.78
EAzY Larson LyaRed 3076 -0.41 4.96 -0.32 4.09 8.91
@@
zphot 1827 0.81 4.23 0.76 3.50 6.79
EAzY v1.3 1825 -0.22 4.31 0.04 3.01 8.82
EAzY Larson Lya 1820 -0.74 4.10 -0.53 2.91 8.52
EAzY Larson LyaRed 1820 -0.78 4.11 -0.54 2.97 8.74
@@
zphot 2016 0.43 4.00 0.34 3.02 5.80
EAzY v1.3 2014 -0.56 4.05 -0.32 2.90 10.03
EAzY Larson Lya 2012 -0.74 3.80 -0.47 2.84 9.94
EAzY Larson LyaRed 2012 -0.84 3.83 -0.54 2.87 10.24
@@
zphot 4298 1.37 4.60 1.35 4.16 7.00
EAzY v1.3 4293 -0.09 4.56 0.06 3.64 7.43
EAzY Larson Lya 4284 -0.68 4.69 -0.41 3.90 9.03
EAzY Larson LyaRed 4284 -0.68 4.70 -0.41 3.92 9.15
@@
zphot 3710 1.03 4.50 0.96 3.80 5.80
EAzY v1.3 3708 -0.06 4.54 0.09 3.65 5.18
EAzY Larson Lya 3702 -0.65 4.85 -0.43 3.85 7.43
EAzY Larson LyaRed 3702 -0.65 4.85 -0.46 3.87 7.59
In Figs. 17 we show the comparison between the spectroscopic and photometric redshifts, the latter computed as the median of the three runs with (i) zphot, (ii) EAzY with eazy_v1.3, and (iii) EAzY with Larson et al. (2023b) FSPS Set 1 + Set 4 templates, as described in Sect. 5. Table 8 summarizes the statistics for all the four runs.
Figure 17: Accuracy of photometric redshifts (median of the three runs described in Sect. 5) on spectroscopic samples, for all fields together (top left panel), and separately for each field. In each panel, N is the number of sources; $f_{outliers}$ is the percentage of sources with $|z_{phot}-z_{spec}|/(1+z_{spec})>0.15$; the other statistics are computed on non-outliers, with NMAD=$1.48\times\mbox{median}[|z_{phot}-z_{spec}|/(1+z_{spec})]$. Color coding is $<\mbox{SNR}>=(\mbox{SNR}_{\mbox{F444W}}+\mbox{SNR}_{\mbox{F356W}}+\mbox{SNR}_{\mbox{F277W}}+\mbox{SNR}_{\mbox{F200W}})/4$.

### v3

@@ 6 Summary and conclusions
We present and discuss a major release of photometric catalogues, collecting data from eight JWST observational programmes on six deep extra-galactic fields. We created a new reduction of the Abell 2744 composite mosaics, gathering the GLASS-JWST, UNCOVER, DDT2756, and GO3990 programmes. For the other fields, we used the reductions provided by the teams of the other programmes. The NIRCam data are complemented with archival HST images, which we used to obtain new measurements.
The catalogues, mainly conceived for high-redshift science, include a grand-total of 531,173 objects, of which 18,563 are tagged as spurious and 2,217 more have bad photometric measurements ($\geq$400; see Sect. 3.5). Sources were detected with SExtractor on stacks of the F356W and F [^4] W mosaics of each field. For all the detections, we provide positions in equatorial and pixel coordinates, basic morphological parameters, total fluxes, and corresponding uncertainties in 16 photometric bands computed by means of the software a-phot, a diagnostic flag, and four photometric redshift estimates obtained with zphot and EAzY.
We performed validation tests on the astrometry, the photometry and the photo-$z$ accuracy, comparing our catalogues against other releases and finding a general good agreement. We encourage the community to exploit these catalogues for any suitable scientific purpose.
7 Data availability
The catalogues are available for download from the Astrodeep website [^16] https://astrodeep.eu/astrodeep-jwst-catalogs/. Updates and/or additional releases will be documented on the website. It is also possible to visualize and explore the catalogues on the Cosmological Surveys Rainbow Database [^17] http://arcoirix.cab.inta-csic.es/
Rainbow_navigator_public/
(Pérez-González et al., 2005, 2008; Barro et al., 2019).
The appendix sections of this paper are available on Zenodo [^18] https://zenodo.org/records/13886578.
The ‘optimal’ photometric catalogues (see Sect. 3.4) and the photo-$z$ catalogues for all fields are also available in electronic form at the CDS via anonymous ftp to cdsarc.u-strasbg.fr (130.79.128.5) or via http://cdsweb.u-strasbg.fr/cgi-bin/qcat?J/A+A/.
@@ Acknowledgements.
We also acknowledge support from the INAF Theory Grant 2022 “Forward modeling of theoretical models for upcoming surveys” (PI Merlin), Large Grant 2022 “Extragalactic Surveys with JWST” (PI Pentericci), Mini Grant 2022 “The evolution of passive galaxies through cosmic time” (PI Santini) and Mini Grant 2022 “Reionization and Fundamental Cosmology with High-Redshift Galaxies” (PI Castellano).
@@
P.G.P.-G. acknowledges support from grant PID2022-139567NB-I00 funded by Spanish Ministerio de Ciencia, Innovación y Universidades MCIU/AEI/10.13039/501100011033, and the European Union through FEDER Una manera de hacer Europa.
Appendix A Catalogue format
We release several catalogues for each field, among which one (named “optap”-catalogue) is obtained assigning to each source the colours computed in its “optimal” aperture (see Section 3.4). The catalogues are published in fits format, and come with a README file that explains the meaning and format of the columns. All catalogues first list a set of columns indicating the unique identifiers of the detected objects, their position in equatorial coordinates and in pixels, and some values from the SExtractor detection run on the F356W+F [^4] W stack mosaics: ISOAREA_IMAGE, CLASS_STAR, FLAGS, FLUX_RADIUS, FLUX_AUTO, FLUXERR_AUTO. Then, some measures obtained with a-phot again on the detection stack are listed: the major semi-axis of the elliptical isophote, the ellipticity and the position angle, and the Kron radius (in pixels). The total fluxes in the 16 photometric bands, computed as described in Section 3.5, are then given, followed by the corresponding 16 uncertainties (all in $\mu$Jy units). Finally, the last column lists the flags discussed in Section 3.5. In the “optap”-catalogue, the value of the “optimal” aperture is also given (in arcseconds). We also release the original raw catalogue containing all the measured aperture fluxes.
We separately release a set of photometric redshift catalogues estimated from the “optap”-catalogue with local background subtraction. For each source we list the ID, RA and Dec, the spectroscopic redshift when available, the four estimates of the photometric redshift (see Section 5), and the flag assigned to the source as reported also in the “optap”-catalogue, with the additional indication of a negative sign for the sources lacking information in the bands necessary to identify the Lyman break.
@@ Appendix B Astrometry
As explained in Section 3.1, we re-aligned HST mosaics to the NIRCam grid by means of an automatic procedure. Table 6 lists the offsets in RA and Dec after the procedure, for all the bands and fields. We point out that the CANDELS EGS and CEERS HDR1 HST data are obtained from the same observations, but the latter has been re-aligned to Gaia-DR3; since we took care to correct all of the images for astrometric accuracy, the two datasets should be perfectly consistent.
Table 6: The median and MAD of the astrometric offsets of the HST bands.
Band $N_{obj.}$ $\Delta{RA}$ (”) $\Delta{DEC}$ (”)
ABELL2744
F435W 1307 0.009 $\pm$ 0.02 $0$$\pm$ 0.01
@@
 [^19] The values are computed by cross-matching $N_{obj}$ sources and corrected before re-projecting the HST images to the JWST grid. For each field, the F160W values are the coordinate offsets between the original HST F160W and the NIRCam F150W band, while the others refer to the offsets between the given band and F160W band after it was re-aligned to the F150W.
Appendix C Selection of point-like and potentially spurious sources
@@
We point out that this approach joins two similar and complementary techniques using diagnostic diagrams to identify the loci occupied by point-like and spurious sources, i.e. the S/N vs radius plane and the $\mu$-mag vs mag plane. Fig. 13 shows the PCA plane for all fields in the left column, and the two diagrams corresponding to the other techniques in the central and right columns.
@@
Figure 13: Star/galaxy separation and flagging of spurious sources. From left to right: as obtained in the PCA plane described in Section 3.5, projected on the hlr vs. S/N plane, projected on the $\mu$MAX-MAG vs. MAG plane. The last column shows the values of the FWHM in the detection F356W+F [^4] W stack, for the sources identified as point-source.
Appendix D Comparisons with other photometric catalogues
Fig. 14 to 19 show how the photometry obtained in this work compares to that from other available catalogues. For each field we consider recent NIRCam-based catalogues and archival HST based data, from CANDELS or Astrodeep. For the latter, we used the $Ks$ and IRAC CH1-2 bands as proxies for the F200W, F356W and F [^4] W JWST bands, applying the following colour corrections, based on Bruzual & Charlot (2003) theoretical templates: $Ks$=F200W-0.06, IRAC1=F356W+0.02, IRAC2=F [^4] W-0.01.
Each panel of the plots shows the comparison for one colour, with the relative difference $\Delta c/c$ (i.e. colour measured in this work minus colour in the reference work, divided by colour in the archival work) as a function of the magnitude of the second band in this work (e.g., F [^4] W magnitude if the colour is F356W-F [^4] W). The blue line is the median of the distribution. We consider sources with S/N$>$5 and flag$<$200. See Section 4 for more details.
Figure 14: Comparison of colours measured in this work and archival catalogues. Shown are relative errors $\Delta c$, i.e. the colours measured in this work minus those in the reference catalogues), vs. the second band magnitude in this work catalogue (e.g. F [^4] W for the F356-F [^4] W colour). Top to bottom: ABELL2744 vs. P23 and NGDEEP vs. Finkelstein’s catalogue (priv. comm.). The number of cross-matched sources after excluding those with S/N $<5$ in any of the two catalogues or flag $\geq 200$ in this work is also given in each panel; the blue line is the median of the distribution.
Figure 15: Same as Fig. 14, for (top to bottom) ABELL2744 vs. Weaver et al. (2024) and CEERS vs. Finkelstein et al. (2023).
Figure 16: Same as Fig. 14, for JADES-GN and JADES-GS vs. the JADES team catalogues (Rieke et al. 2023).
Figure 17: Same as Fig. 14, for ABELL2744 vs. Merlin et al. (2016b) showing individual points rather than a density hexbin plot because of the small number of cross-matched sources, and PRIMER-COSMOS vs. Nayyeri et al. (2017). The HST colours are transformed into JWST colours by means of the corrections described in Appendix D.
Figure 19: Same as Fig. 14, JADES-GS vs. Merlin et al. (2021) and NGDEEP vs. Merlin et al. (2021), the latter showing individual points rather than a density hexbin plot because of the small number of cross-matched sources. The HST colours are transformed into JWST colours by means of the corrections described in Appendix D.
Appendix E Photometric redshift validation
In Fig. 20 we show the comparison between the spectroscopic and photometric redshifts, the latter computed as the median of the three runs with (i) zphot, (ii) EAzY with eazy_v1.3, and (iii) EAzY with Larson et al. (2023b) FSPS Set 1 + Set 4 templates, as described in Section 5. Table 8 summarises the statistics for all four runs.
@@ Appendix E Photometric redshift validation
Table 8: Comparative statistics for photometric vs. spectroscopic redshifts.
@@
zphot 1355 0.57 4.28 0.47 3.33 10.04
EAzY v1.3 1354 -0.81 4.31 -0.46 3.36 13.81
EAzY Larson Lya 1338 -0.15 4.64 -0.28 3.29 13.75
EAzY Larson LyaRed 1338 -0.32 4.70 -0.33 3.46 14.28
@@
zphot 3186 1.17 4.63 0.89 3.64 7.97
EAzY v1.3 3182 0.38 4.55 0.41 3.59 7.54
EAzY Larson Lya 3180 -0.40 4.96 -0.26 4.04 9.03
EAzY Larson LyaRed 3180 -0.41 4.97 -0.30 4.08 9.18
@@
zphot 1848 0.81 4.25 0.75 3.47 6.71
EAzY v1.3 1846 -0.24 4.30 0.01 2.98 8.88
EAzY Larson Lya 1841 -0.74 4.08 -0.53 2.89 8.53
EAzY Larson LyaRed 1841 -0.77 4.10 -0.55 2.93 8.75
@@
zphot 2021 0.41 4.01 0.32 3.02 5.79
EAzY v1.3 2019 -0.57 4.05 -0.35 2.91 10.10
EAzY Larson Lya 2017 -0.75 3.80 -0.47 2.81 9.92
EAzY Larson LyaRed 2017 -0.85 3.83 -0.54 2.84 10.21
@@
zphot 4413 1.37 4.61 1.32 4.15 6.80
EAzY v1.3 4408 -0.03 4.59 0.10 3.64 7.17
EAzY Larson Lya 4399 -0.63 4.69 -0.36 3.85 8.98
EAzY Larson LyaRed 4399 -0.65 4.71 -0.37 3.89 9.09
@@
zphot 3721 1.03 4.50 0.95 3.80 5.80
EAzY v1.3 3719 -0.05 4.55 0.09 3.65 5.22
EAzY Larson Lya 3713 -0.64 4.86 -0.43 3.85 7.41
EAzY Larson LyaRed 3713 -0.65 4.86 -0.46 3.87 7.57
202020 The listed values are: number N of matched sources (the number can vary due to failings in the fitting procedures of the different codes); mean, standard deviation, median and NMAD of the quantity $|z_{phot}-z_{spec}|/(1+z_{spec})$, all multiplied by 100 to make the differences clearer; and percentage of outliers $f_{outliers}$ as described in Section 5, for each run using different software and/or templates, and for each field.